OpenStudy (anonymous):

I know the radius of the whole dart board, but it asks to find the radius of the inner most circle inside the dart board. How can I find it?

ganeshie8 (ganeshie8):

wat other info it gives ?

OpenStudy (anonymous):

there were 350 holes in the dart board that seemed evenly spaced. of these, 5 were within circle 9(the most inner circle). the radius of the whole board is 8 inches. it asks to find a good estimate for the radius of circle 9.

OpenStudy (anonymous):

is this enough information? it's all the problem states

ganeshie8 (ganeshie8):

yes it is enough, one sec..

ganeshie8 (ganeshie8):

Since there are 5 holes in the inner circle, and 350 holes in the bigger circle, ratio of areas would be :- \(\large \frac{5}{350} = \frac{1}{70}\)

ganeshie8 (ganeshie8):

that gives, ratio of radii = \(\large \sqrt{\frac{1}{70}}\)

ganeshie8 (ganeshie8):

Multiply it wid the bigger radius

OpenStudy (anonymous):

thank you so much!

ganeshie8 (ganeshie8):

np :)